OpenStudy (fudgyfingers):

Evaluate the integral of x^2(x^3+1)^3

OpenStudy (bmk614):

You can use U, du substitution where u=x^3+1. Do you know how to do this?

OpenStudy (fudgyfingers):

Not really, I didn't understand it when it was explained to me

OpenStudy (bmk614):

when u=x^3+1 to find du, you take the derivative of u

OpenStudy (bmk614):

What do you get as du?

OpenStudy (fudgyfingers):

3x^2

OpenStudy (bmk614):

du=3x^2dx In order to substitute dx you replace everything you can with du. Since we only have a x^2dx, we have to get rid of the 3. You can get 1/3du= x^2dx. Just replace x^2dx with the 1/3 du, and replace x^3+1 with u

OpenStudy (bmk614):

what do you get?

OpenStudy (fudgyfingers):

I'm confused as to where I should be plugging that in, I'm sorry

OpenStudy (bmk614):

well where you see the x^3+1, take that out and put a u

OpenStudy (fudgyfingers):

from the original integral i typed out?

OpenStudy (bmk614):

Yes.

OpenStudy (fudgyfingers):

But you said u= x^3 +1

OpenStudy (bmk614):

yes, since x^3+1 is the same thing as u, you can just substitute it.

OpenStudy (fudgyfingers):

integral of x^2 u dx

OpenStudy (fudgyfingers):

*u^3

OpenStudy (bmk614):

Yes, but then subsititute the x^2 dx with 1/3du

OpenStudy (fudgyfingers):

okay, so integral of u^3 * 1/3 du

OpenStudy (fudgyfingers):

is that right?

OpenStudy (bmk614):

Yes. Now take the integral of that.

OpenStudy (fudgyfingers):

I realized what I've been doing wrong

OpenStudy (fudgyfingers):

would you be willing to check the answer for me?

OpenStudy (bmk614):

Yes. But what did you get at the integral of that in terms of u

OpenStudy (fudgyfingers):

(u^4)/12 +c, then you just go back and plug in the value for you and you get the answer of (x^3+1)^4/12 + c

OpenStudy (fudgyfingers):

thank you so much!

OpenStudy (bmk614):

No problem! I am glad I could help!